# Pair Production in time-dependent Electric field at Finite times

### v5

@@ III Pair Production in Sauter-pulse electric field
A spatially uniform external background is a common approximation of the electromagnetic field near the focal region of two counter-propagating laser pulses along the z-axis, generating a standing wave Nakajima (2004); Blaschke et al. (2009). In general, the pair-production process takes place close to the electric field maximum ( comparable to the critical field limit), where the magnetic field vanishes. Despite the fact that laser fields typically include many optical cycles, in this case, we examine a relatively simple model of the external field made up of the Sauter profile, which can be thought of as an extremely short laser pulse.
@@
[…]
@@
$\left(\frac{4}{\tau^{2}}y\left(1-y\right)\partial_{y}y\left(1-y\right)\partial_{y}+\omega^{2}(\bm{p},y)+4\textrm{i}E_{0}y\left(1-y\right)\right)\psi_{\bm{p}}(y)=0.$ (53)
@@
$\omega_{0}=\omega(\bm{p},y=0)=\sqrt{\epsilon_{\perp}^{2}+P_{0}^{2}},\quad\omega_{1}=\omega(\bm{p},y=1)=\sqrt{\epsilon_{\perp}^{2}+P_{1}^{2}}.$ (56)
@@
[…]
@@ III.1 Particle distribution function
$|\beta(\bm{p},y)|^{2}=\frac{\epsilon^{2}_{\perp}(p_{\perp})}{2\omega(\bm{p},y)\bigl{(}\omega(\bm{p},y)-P(p_{\parallel},y)\bigr{)}}\biggl{|}\Biggl{(}\frac{2}{\tau}y(1-y)\partial_{y}+\textrm{i}\omega(\bm{p},y)\Biggr{)}\psi^{(+)}(\bm{p},y)\biggr{|}^{2}.$ (71)
@@
where $|N^{(+)}(\bm{p})|^{2}=\frac{1}{{2\omega_{0}\bigl{(}\omega_{0}\mp P_{0}\bigr{)}}},f_{1}={}_{2}\mathcal{F}_{1}\left(1+a,1+b,1+c;y\right)$, and $f_{2}={}_{2}\mathcal{F}_{1}\left(a,b,c;y\right).$
This expression for the time-dependent single-particle distribution function provides insight into pair production at various dynamical stages Blaschke et al. (2011a); Banerjee and Singh (2022). In this context, we focus on analyzing the evolution of the distribution function from the initial QEPP stage to the final REPP stage, passing through the transient stage Banerjee and Singh (2022). Within this framework, the distribution function encodes the spectral information of the produced particles, and its behavior is discussed in Sec. IV. Additionally, in Sec. III.3, we derive a simple approximate expression for the distribution function, valid for $t\gg\tau$.
@@ III.2 Correlation function
Quantum Vacuum in the presence of an external electric field may be described by a complex correlation function Banerjee and Singh (2022, 2019). It is defined as,
@@
In several research articles Schmidt et al. (1998); Banerjee and Singh (2019); Blaschke et al. (2017); Fedotov et al. (2011); Banerjee and Singh (2022), the particle-antiparticle correlation function is redefined by incorporating the slowly varying component of the time-dependent creation and annihilation operators in adiabatic number basis:
^$\hat{\mathcal{B}}_{{\bm{p}}r}(t)=\hat{B}_{{\bm{p}}r}(t)\textrm{e}^{-\textrm{i}\Theta_{\bm{p}}(t_{0},t)},$ (75)
@@
Vacuum polarization effects play a fundamental role in the process of pair production, influencing the dynamics of particle creation in strong fields. These effects are captured by the correlation function $\mathcal{C}(\bm{p},t)$, which can be decomposed into its real and imaginary parts to describe different aspects of the process  Banerjee and Singh (2019); Aleksandrov:2020mez; Smolyansky:2019dqd. Specifically, the real part of $\mathcal{C}(\bm{p},t)$ is denoted as $u(\bm{p},t)=\Re\left(\mathcal{C}(\bm{p},t)\right)$, while the imaginary part is given by $v(\bm{p},t)=\Im\left(\mathcal{C}(\bm{p},t)\right)$. These components are crucial for understanding the underlying mechanisms of pair production and its dependence on time and momentum.
@@
The function $u(\bm{p},t)$ depicts vacuum polarization effects and pair production phenomena. The function $v(\bm{p},t)$ serves as a counterterm to pair production, effectively representing pair annihilation in the vacuum excitation process Aleksandrov:2020mez. A comprehensive discussion of these insights is presented in detail in Sec. IV.3.
@@ III.3 Approximate analytical expression for the distribution function at finite time
First, we start with approximating the Gauss-hypergeometric function as $y\to 1$. It is crucial to ensure smooth convergence towards the limit of ${}_{2}\mathcal{F}_{1}\left(a,b,c;y\rightarrow 1\right)$. This task is complicated by the intricate nature of the parameters $a$, $b$, and $c$ in this specific context, which makes it essential to exercise caution when dealing with this limit. Therefore, it is beneficial to transform the argument by substituting $y$ with $(1-y)$. This transition can be achieved using the following mathematical identity
@@
$f_{2}$ $=\frac{\Gamma(c)\Gamma(c-a-b)}{\Gamma(c-a)\Gamma(c-b)}\Biggl{(}1+\frac{ab}{1+a+b-c}(1-y)+...\Biggr{)}$
$+(1-y)^{(c-a-b)}\frac{\Gamma(c)\Gamma(a+b-c)}{\Gamma(a)\Gamma(b)}\Biggl{(}1+\frac{(c-a)(c-b)}{(1+c-a-b)}(1-y)+...\Biggr{)}.$ (84)
@@
As time progresses, especially in the case with significant time intervals such as $t>\tau$, it becomes evident that $(1-y)$ approaches zero. Consequently, the dominant contribution to the particle distribution function as $y\rightarrow 1$ originates from the zeroth-order term, which is independent of $(1-y).$ As discussed in many literature for deriving the asymptotic expression for the particle distribution function, the Gauss-hypergeometric functions are truncated up to zeroth order onlyHebenstreit et al. (2010); Gavrilov and Gitman (1996).
Building on the earlier discussion (see Sec. III.1), our focus is primarily on understanding the dynamics of the particle distribution function over finite time rather than just its asymptotic trends. We want to examine the distribution function, $f(\bm{p},y)$, in the vicinity of $y\rightarrow 1.$ In that case, it is necessary to incorporate other higher-order terms in the expression of the asymptotic particle distribution function. It means expanding the time-dependent distribution function defined in Eq. (72) as a power series in the variable $(1-y)$ as:
@@
Therefore,
@@
where we have used
@@
$+\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln{(1-y)}}\Biggl{(}\frac{2E_{0}^{2}\tau^{2}\epsilon^{2}_{\perp}}{\omega_{1}^{3}}+\Bigl{(}(\omega_{1}-\omega_{0})-\frac{2E_{0}\tau}{\omega_{1}}P_{1}\Bigr{)}\frac{(c-a)(c-b)}{(1+c-a-b)}\Biggr{)}\Biggr{)}.$ (89)
@@
We also approximate the cross term $\Gamma_{1}\overline{\Gamma_{2}}$, which appears in the expression for the distribution function. The detailed calculation of this term using the approximation for the Gamma function is provided in Appendix A.1. Here, we simply write final expression for the $\Gamma_{1}\overline{\Gamma_{2}}$ :
$\Gamma_{1}\overline{\Gamma_{2}}$ $\approx\left|\Gamma_{1}\overline{\Gamma_{2}}\right|e^{\textrm{i}\varrho},$ (94)
@@
$\left|\Gamma_{1}\overline{\Gamma_{2}}\right|$ $=\frac{\omega_{0}}{2\omega_{1}}\cdot\frac{\exp\left[\frac{\pi\tau}{2}(\omega_{0}-\omega_{1}+2E_{0}\tau)\right]}{\sinh(\pi\tau\omega_{0})\sqrt{1+\omega_{1}^{2}\tau^{2}}}\sqrt{\frac{(\omega_{0}+2E_{0}\tau)^{2}-\omega_{1}^{2}}{\omega_{1}^{2}-(2E_{0}\tau-\omega_{0})^{2}}},$ (95)
@@
$\qquad\qquad\quad+\frac{\textrm{i}}{\omega_{1}}\left(\frac{2E_{0}^{2}\tau^{2}\epsilon_{\perp}^{2}}{\omega_{1}^{2}}(c-a-b-1)-ab(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\right)\Bigg{)}$
@@
×$\qquad\qquad\qquad\times\left(\frac{2y}{\tau}\omega_{1}(c-a+1)(c-b+1)-(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\right)\Bigg{)}\Bigg{]}\Bigg{|}^{2}.$ (97)
To explore the behavior of the distribution function at finite times $(t>\tau)$, we aim to express $f(\bm{p},y)$ in a series involving $(1-y)$, as discussed previously (refer to Eq. (85)). We can then consider truncating higher-order terms to simplify the analysis while still capturing essential features. In this context, we focus exclusively on terms up to order $(1-y)^{2}$, disregarding higher-order terms in Eq. (97). This approach is sufficient to explain the interesting results discussed later.
After performing extensive calculations, we derive a simplified expression for the particle distribution function in terms of the small parameter $(1-y)$, up to second order. A detailed evaluation of $f(\bm{p},y)$ for better clarity is provided in Appendix A.1. This corresponding expression is given by:
@@
$\mathrm{C}_{1}(\bm{p},y)$ $=4y\Omega_{0}|\Gamma_{1}\overline{\Gamma_{2}}|E_{0}\tau(P_{1}-\omega_{1})\cos{(\Upsilon)}+4y|\Gamma_{2}|^{2}(2\Omega_{0}(\omega_{1}^{2}-\omega_{0}\omega_{1}+(\omega_{1}-P_{1})E_{0}\tau)+\Omega_{1}\omega_{1}^{2}y),$ (100)
@@
$+4\Omega_{2}y^{2}\omega_{1}^{2}+\frac{\Omega_{0}}{4\omega_{1}^{2}(1+\tau^{2}\omega_{1}^{2})}\sigma_{0}\Biggr{]}+\frac{|\Gamma_{1}\overline{\Gamma_{2}}|(\cos{(\Upsilon)}\sigma_{1}+\sin{(\Upsilon)}\sigma_{2})}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}.$ (101)
@@
Also note that $\mathrm{C}_{0}(\bm{p},y)$, $\mathrm{C}_{1}(\bm{p},y)$, and $\mathrm{C}_{2}(\bm{p},y)$ are independent of the function of $(1-y)$.
@@ IV.1 Temporal evolution of particle distribution
Quantum vacuum becomes unstable under the action of the external electric field. As a consequence, virtual particle-antiparticle pairs are created in an off-shell mass configuration. These virtual charged particles are accelerated by the electric field to enough energy to become real particles in an on-shell mass configuration. During the action of the external force, pair annihilation processes occur simultaneously with the pair creation, giving rise to a dynamical quasiparticle plasma Blaschke et al. (2015). This results in different states of in and out-vacuums. The most complete description of vacuum pair creation from in-state to out-state is given by the single-particle distribution function.
The time evolution of the single-particle distribution function shows that electron-positron plasma(EPP) excited from vacuum passes through three different temporal stages: the QEPP stage, the transient stage, and the final REPP in the out-state as indicated by the author in Blaschke et al. (2011a); Banerjee and Singh (2022). This is illustrated in Figure 3. The initial stage QEPP and the final REPP stage are separated by the fast oscillation of EPP by the transient stage. The transient stage is considered to begin at time $t_{in}$, where the oscillation of $f(\bm{p},t)$ first reaches the REPP level. The time $t_{out}$, marking the end of the transient stage, is defined as when the average level of the oscillating $f(\bm{p},t)$ hits the REPP level. After this point, the REPP stage begins. At the REPP stage, quasi-particles become independent, and real particle-antiparticles are observed with a lesser value off than that at the electric field maximum at $t=0.$ Each of the three stages contributes to various physical effects, including vacuum polarization effects Smolyansky et al. (2010), the emission of annihilation photons from the focal point of colliding laser beams Blaschke et al. (2006); Blaschke et al. (2011b, c, a), birefringence effects Di Piazza et al. (2006), and other secondary phenomena. To accurately estimate the contributions of these stages to observable effects, a detailed analysis of each period of the EPP’s evolution is quite helpful.
Since the particle distribution function depends on momentum $p_{\parallel}=p_{3}$ parallel to the direction of the electric field and the modulus of the transverse momentum $p_{\perp}=\sqrt{p_{1}^{2}+p_{2}^{2}}.$ Due to the rotation symmetry of the problem about the field axis, one may parameterize the momentum vector as $\bm{p}=(p_{1},0,p_{3})$ with the transverse component, $p_{\perp}=p_{1}$, and longitudinal component $p_{\parallel}=p_{3}.$ Furthermore, the quasienergy $\omega(\bm{p},t)$, the transverse energy ϵ(p)subscriptitalic-$\epsilon_{\perp}(p_{\perp})$ and the longitudinal quasi-momentum $P(p_{\parallel},t)$ are defined as
@@
ϵ(p)subscriptitalic-$\epsilon_{\perp}(p_{\perp})$ $=\sqrt{m^{2}+p^{2}_{\perp}},$ (107)
@@
Figure 4: The relationship between momentum and the transient domain’s time of occurrence. Left panel: for longitudinal momentum.Right panel: for transverse momentum. The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}]$ All units are in electron mass units.
          $p_{\parallel}[m]$           $t_{in}[m^{-1}]$           $f(p_{\parallel},t_{in})$           $t_{out}[m^{-1}]$           $f(p_{\parallel},t_{out})$
          0.0           12.57           $7.203\times 10^{-7}$           24.038           $3.309\times 10^{-7}$
          0.25           13.37           $5.576\times 10^{-7}$           24.803           $3.112\times 10^{-7}$
          0.50           13.78           $2.242\times 10^{-7}$           24.320           $1.599\times 10^{-7}$
          0.75           14.53           $7.769\times 10^{-8}$           25.799           $7.033\times 10^{-8}$
          1.00           16.76           $1.748\times 10^{-8}$           26.866           $1.082\times 10^{-8}$
Table 1: Transient region time labeled by longitudinal momentum $p_{\parallel}.$The time interval during which a strong oscillation emerges in the evolution of the distribution function is referred to as $t_{in}$, and $t_{out}$ marks the point where this vigorous oscillation diminishes, leading to a residual distribution function. The value of transverse momentum is taken to be zero, and all the units are taken in electron mass unit. The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}].$
Next, we examine the influence of transverse momentum on temporal stages, as depicted in the right panel of Fig. 4. We observe intriguing behavior where the distribution function slowly approaches the REPP stage with a residual value $f(p_{\parallel}=0,p_{\perp},t>t_{out})$, which is lowest compared to small transverse momentum values $p_{\perp}$. The behavior of these stages is primarily determined by the double quasi-energy $2\omega(\bm{p},t)$. Increasing the transverse momentum value, and thus the transverse energy ϵ(p)subscriptitalic-$\epsilon_{\perp}(p_{\perp})$, increases the dynamical energy gap $2\omega(\bm{p},t)$. Consequently, achieving the on-shell mass configuration takes a longer time, see the right panel of Fig. 4.

### v6

@@ III Pair Production in Sauter-pulse electric field
A spatially uniform external field background is a common approximation of the electromagnetic field near the focal region of two counter-propagating laser pulses along the z-axis, generating a standing wave Nakajima (2004); Blaschke et al. (2009). In general, the pair-production process takes place close to the electric field maximum (comparable to the critical field limit), where the magnetic field vanishes. Despite the fact that laser fields typically include many optical cycles, in this case, we examine a relatively simple model of the external field made up of the Sauter profile, which can be thought of as an extremely short laser pulse.
@@
where $\tau$ is the duration of pulse, and $E_{0}$ is field strength. This electric field exponentially goes to zero for $|t|>>\tau.$ In the limit of $\tau\rightarrow\infty$ the electric field becomes homogeneous in time. We can choose a gauge in which $A_{0}=0$ and the vector potential associated with the electric field is $(0,0,A(t)=-\int dtE(t)).$ After the integration, we find the Sauter-type gauge potential,
@@
$\left[\frac{4}{\tau^{2}}y\left(1-y\right)\partial_{y}y\left(1-y\right)\partial_{y}+\omega^{2}(\bm{p},y)+4\textrm{i}E_{0}y\left(1-y\right)\right]\psi_{\bm{p}}(y)=0.$ (53)
@@
$\omega_{0}=\omega(\bm{p},y=0)=\sqrt{\epsilon_{\perp}^{2}(p_{\perp})+P_{0}^{2}},\quad\omega_{1}=\omega(\bm{p},y=1)=\sqrt{\epsilon_{\perp}^{2}(p_{\perp})+P_{1}^{2}}.$ (56)
@@
$\Theta_{\bm{p}}(y_{0},y)=\frac{\tau}{2}\int^{y}_{y_{0}}dy^{\prime}\frac{\omega(\bm{p},y^{\prime})}{y^{\prime}\left(1-y^{\prime}\right)},$ (67)
@@ III.1 Particle distribution function
$|\beta_{\bm{p}}(y)|^{2}=\frac{\epsilon^{2}_{\perp}(p_{\perp})}{2\omega(\bm{p},y)\bigl(\omega(\bm{p},y)-P(p_{\parallel},y)\bigr)}\left|\Biggl(\frac{2}{\tau}y(1-y)\partial_{y}+\textrm{i}\omega(\bm{p},y)\Biggr)\psi^{(+)}_{\bm{p}}(y)\right|^{2}.$ (71)
@@
where $|N^{(+)}(\bm{p})|^{2}=\frac{1}{{2\omega_{0}\bigl(\omega_{0}-P_{0}\bigr)}},f_{1}={}_{2}\mathcal{F}_{1}\left(1+a,1+b,1+c;y\right)$, and $f_{2}={}_{2}\mathcal{F}_{1}\left(a,b,c;y\right).$
This expression for the time-dependent single-particle distribution function provides insight into pair production at various dynamical stages. In this context, we focus on analyzing the evolution of the distribution function from the initial QEPP stage to the final REPP stage, passing through the transient stage Banerjee and Singh (2022). Within this framework, the distribution function encodes the spectral information of the produced particles, and its behavior is discussed in Sec. IV. Additionally, in Sec. III.3, we derive a simple approximate expression for the distribution function, valid for $t\gg\tau$.
@@ III.2 Correlation function
The quantum vacuum in the presence of an external electric field may be described by a complex correlation function. It is defined as,
@@
In several research articles Schmidt et al. (1998); Blaschke et al. (2017); Fedotov et al. (2011), the particle-antiparticle correlation function is redefined by incorporating the slowly varying component of the time-dependent creation and annihilation operators in adiabatic number basis:
$\hat{\mathcal{B}}_{{\bm{p}}r}(t)=\hat{B}_{{\bm{p}}r}(t)\textrm{e}^{-\textrm{i}\Theta_{\bm{p}}(t_{0},t)},$ (75)
@@
Vacuum polarization effects play a fundamental role in the process of pair production, influencing the dynamics of particle creation in strong fields. These effects are captured by the correlation function $\mathcal{C}(\bm{p},t)$, which can be decomposed into its real and imaginary parts to describe different aspects of the process Banerjee and Singh (2019); Smolyansky et al. (2019). Specifically, the real part of $\mathcal{C}(\bm{p},t)$ is denoted as $u(\bm{p},t)=\Re\left(\mathcal{C}(\bm{p},t)\right)$, while the imaginary part is given by $v(\bm{p},t)=\Im\left(\mathcal{C}(\bm{p},t)\right)$. These components are crucial for understanding the underlying mechanisms of pair production and its dependence on time and momentum.
@@
The function $u(\bm{p},t)$ depicts vacuum polarization effects and pair production phenomena. The function $v(\bm{p},t)$ serves as a counterterm to pair production, effectively representing pair annihilation in the vacuum excitation process. A comprehensive discussion of these insights is presented in detail in Sec. IV.3.
@@ III.3 Approximate analytical expression for the distribution function at finite time
First, we start with approximating the Gauss-hypergeometric function as $y\to 1$. It is crucial to ensure smooth convergence towards the limit of $_{2}\mathcal{F}_{1}\left(a,b,c;y\rightarrow 1\right)$. This task is complicated by the intricate nature of the parameters $a$, $b$, and $c$ in this specific context, which makes it essential to exercise caution when dealing with this limit. Therefore, it is beneficial to transform the argument by substituting $y$ with (1y)(1-y). This transition can be achieved using the following mathematical identity
@@
[…]
@@
As time progresses, especially in the case with significant time intervals such as $t>\tau$, it becomes evident that (1y)(1-y) approaches zero. Consequently, the dominant contribution to the particle distribution function as $y\rightarrow 1$ originates from the zeroth-order term, which is independent of (1y).(1-y). As discussed in many literature for deriving the asymptotic expression for the particle distribution function, the Gauss-hypergeometric functions are truncated up to zeroth order onlyHebenstreit et al. (2010); Gavrilov and Gitman (1996).
Building on the earlier discussion (see Sec. III.1), our focus is primarily on understanding the dynamics of the particle distribution function over finite time rather than just its asymptotic trends. We want to examine the distribution function, $f(\bm{p},y)$, in the vicinity of $y\rightarrow 1.$ In that case, it is necessary to incorporate other higher-order terms in the expression of the asymptotic particle distribution function. It means expanding the time-dependent distribution function defined in Eq. (72) as a power series in the variable (1y)(1-y) as:
@@
Let us begin by approximating the various terms appearing in the particle distribution functions Eqs. (83) and (84), retaining contributions up to second order and neglecting higher-order terms. Accordingly, we obtain
@@
where we have defined
@@
Here, $\Gamma(\cdot)$ denotes the Gamma function, while $\Gamma_{1}$ and $\Gamma_{2}$ are shorthand coefficients constructed from ratios of Gamma functions.
@@
$+(1-y)^{2}\Biggl(\Gamma_{1}(c-a-b-1)\Biggl(\frac{2E_{0}^{2}\tau^{2}\epsilon_{\perp}^{2}(p_{\perp})}{\omega_{1}^{3}}+\frac{ab(\omega_{1}(\omega_{1}-\omega_{0})-2P_{1}E_{0}\tau)}{\omega_{1}(1+a+b-c)}\Biggl)$
$+\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln{(1-y)}}\Biggl(\frac{2E_{0}^{2}\tau^{2}\epsilon_{\perp}^{2}(p_{\perp})}{\omega_{1}^{3}}+\Bigl((\omega_{1}-\omega_{0})-\frac{2E_{0}\tau}{\omega_{1}}P_{1}\Bigr)\frac{(c-a)(c-b)}{(1+c-a-b)}\Biggr)\Biggr).$ (89)
@@
We also approximate the cross term $\Gamma_{1}\Gamma^{*}_{2}$, which appears in the expression for the distribution function. The detailed calculation of this term using the approximation for the Gamma function is provided in Appendix A. Here, we simply write final expression for the $\Gamma_{1}\Gamma^{*}_{2}$ :
$\Gamma_{1}\Gamma^{*}_{2}$ $\approx\left|\Gamma_{1}\Gamma^{*}_{2}\right|e^{\textrm{i}\varrho},$ (94)
@@
$\left|\Gamma_{1}\Gamma^{*}_{2}\right|$ $=\frac{\omega_{0}}{2\omega_{1}}\cdot\frac{\exp\left[\frac{\pi\tau}{2}(\omega_{0}-\omega_{1}+2E_{0}\tau)\right]}{\sinh(\pi\tau\omega_{0})\sqrt{1+\omega_{1}^{2}\tau^{2}}}\sqrt{\frac{(\omega_{0}+2E_{0}\tau)^{2}-\omega_{1}^{2}}{\omega_{1}^{2}-(2E_{0}\tau-\omega_{0})^{2}}},$ (95)
@@
$\qquad\qquad\quad+\frac{\textrm{i}}{\omega_{1}}\left(\frac{2E_{0}^{2}\tau^{2}\epsilon_{\perp}^{2}(p_{\perp})}{\omega_{1}^{2}}(c-a-b-1)-ab(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\right)\Bigg)$
@@
$\qquad\qquad\qquad\times\left(\frac{2y}{\tau}\omega_{1}(c-a+1)(c-b+1)-(\omega_{1}^{2}-\omega_{1}\omega_{0}-2E_{0}\tau P_{1})\right)\Bigg)\Bigg]\Bigg|^{2}.$ (97)
To explore the behavior of the distribution function at finite times $(t>\tau)$, we aim to express $f(\bm{p},y)$ in a series involving (1y)(1-y), as discussed previously (refer to Eq. (85)). We can then consider truncating higher-order terms to simplify the analysis while still capturing essential features. In this context, we focus exclusively on terms up to order $(1-y)^{2}$, disregarding higher-order terms in Eq. (97). This approach is sufficient to explain the interesting results discussed later.
After performing extensive calculations, we derive a simplified expression for the particle distribution function in terms of the small parameter (1y)(1-y), up to second order. A detailed evaluation of $f(\bm{p},y)$ for better clarity is provided in Appendix A. This corresponding expression is given by:
@@
$\mathrm{C}_{1}(\bm{p},y)$ $=4y\Omega_{0}|\Gamma_{1}\Gamma^{*}_{2}|E_{0}\tau(P_{1}-\omega_{1})\cos{(\Upsilon)}+4y|\Gamma_{2}|^{2}(2\Omega_{0}(\omega_{1}^{2}-\omega_{0}\omega_{1}+(\omega_{1}-P_{1})E_{0}\tau)+\Omega_{1}\omega_{1}^{2}y),$ (100)
@@
$+4\Omega_{2}y^{2}\omega_{1}^{2}+\frac{\Omega_{0}}{4\omega_{1}^{2}(1+\tau^{2}\omega_{1}^{2})}\sigma_{0}\Biggr]+\frac{|\Gamma_{1}\Gamma^{*}_{2}|(\cos{(\Upsilon)}\sigma_{1}+\sin{(\Upsilon)}\sigma_{2})}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}.$ (101)
@@
Also note that $\mathrm{C}_{0}(\bm{p},y)$, $\mathrm{C}_{1}(\bm{p},y)$, and $\mathrm{C}_{2}(\bm{p},y)$ are independent of the function of (1y)(1-y).
@@ IV.1 Temporal evolution of particle distribution
Quantum vacuum becomes unstable under the action of the external electric field. As a consequence, virtual particle-antiparticle pairs are created in an off-shell configuration. These virtual charged particles are accelerated by the electric field to enough energy to become real particles in an on-shell mass configuration. During the action of the external force, pair annihilation processes occur simultaneously with the pair creation, giving rise to a dynamical quasiparticle plasma Blaschke et al. (2015). This results in different states of in and out-vacuums. The most complete description of vacuum pair creation from in-state to out-state is given by the single-particle distribution function.
The time evolution of the single-particle distribution function shows that electron-positron plasma(EPP) excited from vacuum passes through three different temporal stages: the QEPP stage, the transient stage, and the final REPP in the out-state as indicated by the author in Blaschke et al. (2011a); Banerjee and Singh (2022). This is illustrated in Figure 3. The initial stage QEPP and the final REPP stage are separated by the fast oscillation of EPP by the transient stage. The transient stage is considered to begin at time $t_{in}$, where the oscillation of $f(\bm{p},t)$ first reaches the REPP level. The time $t_{out}$, marking the end of the transient stage, is defined as when the average level of the oscillating $f(\bm{p},t)$ hits the REPP level. After this point, the REPP stage begins. At the REPP stage, quasi-particles become independent, and real particle-antiparticles are observed with a lesser value off than that at the electric field maximum at $t=0.$ Each of the three stages contributes to various physical effects, including vacuum polarization effects Smolyansky et al. (2010), the emission of annihilation photons from the focal point of colliding laser beams Blaschke et al. (2006, 2011c, 2011b, 2011a), birefringence effects Di Piazza et al. (2006), and other secondary phenomena. To accurately estimate the contributions of these stages to observable effects, a detailed analysis of each period of the EPP’s evolution is quite helpful.
Since the particle distribution function depends on momentum $p_{\parallel}=p_{3}$ parallel to the direction of the electric field and the modulus of the transverse momentum $p_{\perp}=\sqrt{p_{1}^{2}+p_{2}^{2}}.$ Due to the rotation symmetry of the problem about the field axis, one may parameterize the momentum vector as $\bm{p}=(p_{1},0,p_{3})$ with the transverse component, $p_{\perp}=p_{1}$, and longitudinal component $p_{\parallel}=p_{3}.$ Furthermore, the quasienergy $\omega(\bm{p},t)$, the transverse energy $\epsilon_{\perp}(p_{\perp})$ and the longitudinal quasi-momentum $P(p_{\parallel},t)$ are defined as
@@
$\epsilon_{\perp}(p_{\perp})$ $=\sqrt{m^{2}+p^{2}_{\perp}},$ (107)
@@
Figure 4: The relationship between momentum and the transient domain’s time of occurrence. Left panel: for longitudinal momentum. Right panel: for transverse momentum. The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}]$ . All units are in electron mass units.
   $p_{\parallel}\penalty 10000\ [m]$    $t_{\text{in}}\penalty 10000\ [m^{-1}]$    $f(p_{\parallel},t_{\text{in}})\penalty 10000\ (\times 10^{-7})$    $t_{\text{out}}\penalty 10000\ [m^{-1}]$    $f(p_{\parallel},t_{\text{out}})\penalty 10000\ (\times 10^{-7})$    0.00    12.57    7.203    24.038    3.309    0.25    13.37    5.576    24.803    3.112    0.50    13.78    2.242    24.320    1.599    0.75    14.53    0.776    25.799    0.7033    1.00    16.76    0.174    26.866    0.1082
Table 1: Transient region times labeled by longitudinal momentum $p_{\parallel}$.The transverse momentum is taken to be zero, and all units are in electron mass. The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10$ $m^{-1}$].
The time interval during which a strong oscillation emerges in the evolution of the distribution function is referred to as $t_{in}$, and $t_{out}$ marks the point where this vigorous oscillation diminishes, leading to a residual distribution function.
Next, we examine the influence of transverse momentum on temporal stages, as depicted in the right panel of Fig. 4. We observe intriguing behavior where the distribution function slowly approaches the REPP stage with a residual value $f(p_{\parallel}=0,p_{\perp},t>t_{out})$, which is lowest compared to small transverse momentum values $p_{\perp}$. The behavior of these stages is primarily determined by the double quasi-energy $2\omega(\bm{p},t)$. Increasing the transverse momentum value, and thus the transverse energy $\epsilon_{\perp}(p_{\perp})$, increases the dynamical energy gap $2\omega(\bm{p},t)$. Consequently, achieving the on-shell mass configuration takes a longer time, see the right panel of Fig. 4.